OpenStudy (anonymous):

Input in standard form the equation of the given line. The line that passes through (1, 5) and (-2, 3)

OpenStudy (lncognlto):

To find the equation for the line, we substitute (1, 5) and (-2, 3) into (y - y1)/(x - x1) = (y2 - y1)/(x2 - x1). Therefore y - 5/x - 1 = 3 - 5/-2 - 1 y - 5/x - 1 = -2/-3 y - 5 = (2/3)(x-1) y = 2/3(x) - 2/3 + 5 y = (2/3)(x) -2/3 + 15/3 y = 2/3(x) + 13/3 3y = 2x + 13

OpenStudy (anonymous):

Thank you sooo much!

OpenStudy (lncognlto):

You're welcome. :)